The following information summarizes the standard cost for producing one metal tennis racket frame at Spaulding Industries. In addition, the variances for one month's production are given. Assume that all inventory accounts have zero balances at the beginning of the month.

Standard Cost Per Unit Standard Monthly Costs
Materials $ 5.00 $ 11,000
Direct Labor 3 hrs @ $2.00 6.00 13,200
Factory Overhead:
Variable 2.00 4,400
Fixed 5.00 11,000
$ 18.00 $ 39,600
Variances:
Material price $ 337.50 unfavorable
Material quantity $ 500.00 unfavorable
Labor rate $ 550.00 favorable
Labor efficiency $ 2,150.00 unfavorable

What were the actual direct labor hours worked during the month?

  • 7,675

  • 7,400

  • 6,600

  • 6,325
Answer #1

Answer: Actual Labor hours 7,675

Total Numer of Units Prodcued = 39,600/ 18.00 = 2,200 units

Standard Hours = 2,200 *3 = 6,600

actula Hours

Labour Efficiency Varaince =( Actula Hours - Standard Hours )* Standard Rate

=( Actula Hours -6,600) *2.00 =2,150 = 7,675 (Answer)